June 27, 2019
Left: Maria on the roof of the logging cabin making sure the wall between the drill trench and the core buffer is tight. Right: Steff, Aslak and Chris testing out the rifles in the hills of EGRIP.
The 3 drill shifts are now fully operational operating around the clock, and producing 6 runs a day plus doing the necessary filtering. Still the winch needs gentle operation especially during the ascend of the drill, but it is a pleasure for the loggers to be back at work logging high quality ice cores – even though anti-talk rotation gave a few short runs today.
When the core has been logged, it is stored in the core buffer until processing. Yesterday evening we found out that temperature had been rising too much in the buffer, so today a wall around the logging cabin was re-established to stop circulation of warm air from the drill trench to the core buffer. Buffer temperature is now -19°C (down from -18°C yesterday evening).
On the surface the atmospheric measurements are continuing from the site on the boundary of the clean snow zone, and in the science trench both processing and all measurements are making good progress.
Finally we ended the day in the carpenters garage where Eliza kindly invited everybody for a house-warming get-together after taking over Karl-Emil’s nice spot on the “balcony”.
What we did during the day:
EGRIP population is 33.
Weather: Wind between 5kts and 12kts first from E turning NNW and finally back to NE in the evening. Mostly cloudy with outbreaks of snow, a sunny spell right before dinner. Temperatures from -13°C to -10°C.
FL, Bo Vinther